OpenStudy (anonymous):

Help! Let f(x)= x^2-2x-3. Evaluate and simplify the following difference quotient. f(2+h)-f(2)/h=?

OpenStudy (anonymous):

It says evaluate, so plug those into the formula. f(2) = (2)^2-2(2)-3 f(2+h) = (2+h)^2-2(2+h)-3 Expand and simplify those.

OpenStudy (anonymous):

-3 h^2+2h-3 Where do I go from there?

OpenStudy (anonymous):

Sorry not sure on the next step

OpenStudy (anonymous):

h^2+2h-6?

OpenStudy (anonymous):

You plug those values back into the original one. [f(2+h)-f(2)]/h [h^2+2h-3 - -3]/h [h^2+2h]/h

OpenStudy (anonymous):

If you cancel h in the numerator and denominator you should get h+2 as your answer.

OpenStudy (anonymous):

Got it! thank you for explaining!